Status Report

NASA Space Station On-Orbit Status 31 January 2006

By SpaceRef Editor
February 1, 2006
Filed under , , ,
NASA Space Station On-Orbit Status 31 January 2006

  SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2006) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

Onboard sleep cycle continues to be shifted a total of 7 hours to the right (wake up 8:00am EST, sleep at 11:30pm), to prepare the crew for this week’s EVA-15 spacewalk (which begins with hatch opening at ~5:20pm EST).

Crew activities today were/are mostly centered on the suited dry run preparatory to next Friday’s (2/3) Orlan spacewalk.

After yesterday’s completion of all EVA suit preparations, dry-run activities began today at ~1:10pm EST with CDR McArthur tearing down and removing the air ducts between the Service Module Transfer Compartment (SM PkhO) and DC1 docking compartment/airlock, skipping ventilation fan V3, to make room for the subsequent suited exercise.

At the same time, FE Tokarev began to configure the communications system for the exercise.   [The suited run requires wireless Tranzit-B suit radio telemetry on both semisets and temporary deactivation of the Russian VHF channel 1 (Very High Frequency, Russian: UKV1, for ultra-shortwave) to avoid interference from extraneous radio stations to the Orlans while over Russian ground stations (RGS).  All EVA preps are monitored by the ground via audio.  Tranzit-B TM is to be turned off again tonight art ~5:52pm.]

After another functionality and leak checking of the Orlan-Ms, their equipment and their interface units (BSS) in the DC1 and PkhO, the crew began donning EVA gear at ~2:00pm, including putting on personal gear bags, biomed harness, thermal underwear, LCG (liquid cooling garment), low-noise headset, gloves, etc. 

After another checkout of comm hookups & biomedical parameter telemetry via the BSS interface system for vital signs and equipment monitoring, suiting up will then (~3:25pm) culminate in ingress in the Orlans through their “backdoors” and sealing off of the backpacks.

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV


Next in line are functionality checkouts of the suits and their BSS controls (e.g., temperature control handling, water cooling system ops), preliminary dimensional fit checks at reduced suit pressure (0.4 at, 5.9 psi), and half an hour of testing/training of suited mobility and translation inside the DC1, beginning at ~4:05pm   [These exercises include translation to all DC1 work stations with mated fluid umbilical, verification of Orlan fit, checkout of onboard cooling system operation, assessment of how the interior DC1 config impacts operations with various gear and accessories such as the POV (EVA support panel) and BSS (Orlan interface unit), evaluation of stowage of hardware to be taken out during the spacewalk, plus some typical EVA-15 tasks, such as Russian-American-Russian safety tether swap, Strela crane adapter transfer, Biorisk-2 canister retrieval, etc.]

Egress from the Orlans is timelined for around 4:35pm, to be followed by restoration of communications settings to nominal operation, one hour lunch break, and a two-hour period of post-training cleanup activities (changing clothes, drying out LCG, biomed harness belt, thermal undergarment, socks, comfort gloves, hygienic trunks and comm caps, remove LiOH canister and moisture collector, etc.), and air duct assembly.

Subsequently, after the Orlans are confirmed to be dry, they are to be re-equipped with fresh consumables/replaceable elements for the spacewalk on Friday.

Before beginning the dry run, Valery Tokarev performed the daily routine maintenance of the SM’s environment control & life support system (SOZh), including the toilet system (ASU), and the weekly inspection of the air/liquid condensate separator apparatus (BRPK).

The crew also performed their daily physical exercise program on TVIS treadmill (aerobic, both) and RED resistive exerciser (anaerobic, Bill), today at a reduced 1.5 hrs due to their subsequent strenuous physical activity in the DC1.  Valery’s workout continued his prescribed four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 2 of a new set).

After completion of the suited dry run, the CDR will perform the daily atmospheric status checks for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the CSA-O2 (Compound Specific Analyzer-Oxygen sensor) and CDMK (CO2 Monitoring Kit).  [Last Friday evening (1/27) the first daily CDMK and CSA-O2 readings in the Lab were taken, measuring at 169.3 mmHg and 3.91 mmHg respectively.  These values are within nominal range.  MCC-Houston is requesting these data points in order to better understand some telemetry trends in ppCO2 and ppO2 seen on the ground. Daily status checks are continuing.]


ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

Also later today, the FE will check the operation of the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at 20 degC, including a temperature check on its ART (automatic temperature recorder).   [This daily monitoring/temp checking, carried on the Russian voluntary “time available” task list, will continue until 4/30.]

Yesterday’s maneuver from LVLH YVV (local vertical local horizontal/y-axis in velocity vector) to LVLH +XVV (+x-axis in velocity vector) using the U.S. Thruster Only (USTO) controller went nominal. [The plan is to use USTO controllers for future attitude maneuvers that are not related to visiting vehicles (Soyuz, Progress, etc.)]

Today’s single CEO (Crew Earth Observation) photo target, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, was Pinacates Biosphere Site, New Mexico (the crew had an opportunity to capture winter imagery of this Biosphere Reserve.  High resolution mapping photography will be useful for comparison with summer images).

Over 177,000 of CEO (Crew Earth Observation) images have been taken in the first five years of the ISS.

  CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 12 crew visit:

Expedition 12 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Events Ahead (all dates Eastern; tentative):

  • 02/03/06 — Russian EVA-15
  • 02/11/06 — ISS Reboost Test (in MMOD avoidance mode)
  • 03/03/06 — Progress M-54/19P undocking & reentry
  • 03/30/06 — Soyuz TMA-8/12S launch (Exp. 13 + Marcos Pontes/Brazil)
  • 04/01/06 — Soyuz TMA-8/12S docking (DC1)
  • 04/09/06 — Soyuz TMA-7/11S undocking & reentry
  • 04/16/06 — Soyuz TMA-8/12S relocation (DC1 to FGB nadir port)
  • 04/24/06 — Progress M-56/21P launch
  • 04/26/06 — Progress M-56/21P docking
  • 06/19/06 — Progress M-55/20P undocking & reentry
  • 06/28/06 — Progress M-57/22P launch
  • 06/30/06 — Progress M-57/22P docking.

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.